OpenStudy (anonymous):

convert 1.8 radians into degrees 1. 112.56 2. 103.13 3. 120.65 4. 98.50

OpenStudy (anonymous):

103.132403 degrees

OpenStudy (gorv):

pi radian=180 degree 1 radian=180/pi 1.8 radian=(180/pi)*1.8

OpenStudy (gorv):

103.132

OpenStudy (anonymous):

thanks